OpenStudy (anonymous):

given the reaction 2SO2+O2=2SO3 if 4 moles of SO2 are used then how many moles of SO3 are made? 4moles 2moles 1mole 1/2 mole

OpenStudy (aaronq):

if you look at the coefficients, they're both the same, 2, therefore the moles used and produced will be equivalent. If you wanna use math to show this you can build a relationship between them using proportions/ratios. For a general reaction, aA + bB <-> cC where a,b and c are the stoichiometric coefficients and A,B and C are the species, \[\frac{ Moles of A }{ a }=\frac{ Moles of B }{ b }=\frac{ Moles of C }{ c }\] and so you can plug in your knowns, and solve (algebraically) for your unknown.

OpenStudy (anonymous):

4 moles.